OpenStudy (anonymous):

abody is thrown vertically up from a wall of height 2.2m with a speed of 10m/s the distance travelled in the last second of its motion is ?

OpenStudy (anonymous):

first :) the height doesn't even matter I think. hold on i check

OpenStudy (anonymous):

I think it is .5102040816

OpenStudy (anonymous):

the ans=7m

OpenStudy (anonymous):

crap i read it wrong lol

OpenStudy (aravindg):

do u knw the formula for distance travelled in nth second?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dn=u+a(n-1/2)

OpenStudy (aravindg):

then use that!!

OpenStudy (anonymous):

first we have to find the total tme

OpenStudy (aravindg):

wait a sec that n in equation is nth second ryt?

OpenStudy (anonymous):

s

OpenStudy (aravindg):

k then find the total time

OpenStudy (anonymous):

ok

OpenStudy (aravindg):

take n=total time

OpenStudy (anonymous):

then

OpenStudy (aravindg):

plug that into the eqn

OpenStudy (aravindg):

leme see if i get answer as 7

OpenStudy (anonymous):

use the formula \[s=ut + 0.5at^2\] s= -2.2, u =10, a=-10(assumed) \[-2.2 = 10t -5t^2\]\[5t^2-10t-2.2=0\] solve the equation and you will get time for the whole journey t=2.2s. distance travelled for the last second i think should mean that 1.2 - 2.2 s To find the highest point it reaches, use \[v^2=u^2 + 2as\] where v is the velocity when the stone reaches the highest point which is actually = 0, u=10, a=-10. sub all the values and you will get s=5m. Now to find where does the body reach after 1.2s \[s=ut + 0.5at^2\] u=10,t=1.2,a=-10 sub all the values into the formula you'll get s=4.8m So the distance travelled during the last second= (5-4.8) + 5 + 2.2 =7.4m answer may vary as i used a=-10

OpenStudy (anonymous):

the ans=7m

OpenStudy (anonymous):

u have to take (g=10m/s) it is there in the question

OpenStudy (anonymous):

(5-4.8) + 5 + 2.2=7.8

OpenStudy (anonymous):

I think that you mean the displacement travelled not the distance travelled. then it should be 4.8 + 2.2 = 7

OpenStudy (anonymous):

no its the distance travelled in last second

OpenStudy (anonymous):

distance travelled=7.4 displacement=7 tats the answer i found.

OpenStudy (anonymous):

ok thanzzz